# Accelerating Containerized Service Delivery at the Network Edge

### v1

@@ III-C1 Request Dispatcher
TABLE II: Size distribution of popular image layers. Each percentage represents the proportion of layers that are smaller than the specified threshold.
Thres. (B) % $<$ Thres. Thres. (B) % $<$ Thres.
@@
256 Mi 97.19 605.73 Mi 100
Partial P2P Downloading. We analyzed the layer sizes of the top 100 popular images from Docker Hub as of July 2024, and as shown in Table II, the majority of these layers are relatively small, with a median size of just 1.03 MiB. Based on this observation, Request Dispatcher employs a partial P2P-based downloading strategy that balances efficiency and resource utilization by avoiding the lengthy processes typically associated with full DHT and tracker-based peer discovery [42]. Comprehensive discovery processes can take disproportionately longer than downloading directly from the central registry. However, relying exclusively on registry downloads for these smaller layers could lead to registry overload. To address this, our partial P2P method prioritizes multicast-based local peer discovery, bypassing the more time-consuming DHT and tracker-based methods. This adjustment significantly reduces both data transfer time and volume. If the requested layer is not found on the local network, Request Dispatcher defaults to direct downloads from Docker Hub. Any subsequent requests for the same layer within the local network are then fulfilled internally, ensuring rapid access while minimizing network congestion on the center registry. This approach effectively optimizes the trade-off between fast access and reduced registry load.
III-C2 Popularity- and Network-Aware Downloading
In PeerSync’s P2P downloading process, content is segmented into blocks, a standard practice that enables simultaneous downloading of different blocks from multiple peers, thereby maximizing bandwidth utilization [46]. When Request Dispatcher opts to download a particular image layer, it provides P2P Downloader with the layer details and relevant peer information, initiating the download process. Fig. 4 illustrates the P2P Downloader’s workflow, which consists of five key stages:
  1. 1.
    The process begins with the selection of a batch of blocks for the current download cycle.
  2. 2.
    Potential peers are sampled based on their dynamically updated scores, maintained by Peer Aggregator. This scoring system assesses peers’ suitability for the download task by considering both content popularity and network quality.
  3. 3.
    The peer with the highest score for each block is then selected to fulfill the download request, ensuring the optimal peer is chosen for each block.
  4. 4.
    A download request is sent to the selected peer.
  5. 5.
    Upon receiving a block, its integrity is verified. If the block fails validation, it is returned to the pending queue for re-download; otherwise, it is stored in the cache and delivered to the client.
@@ III-C2 Popularity- and Network-Aware Downloading
The cycle repeats with a new batch of blocks once the current one is successfully downloaded and validated.
@@
Determining the Block Size. The methodology for determining block size and selecting peers for download requests is crucial to optimizing performance. The block size must be carefully balanced: blocks that are too large reduce parallelism by decreasing the number of blocks available for simultaneous download, while blocks that are too small increase protocol overhead and computational load due to a larger Merkle tree and more hash operations. Our experimental analysis, shown in Table III, used an 8194.5 MiB image in a 10 Gbps LAN environment. The results indicated that smaller block sizes significantly prolonged the download time. Based on these findings, we empirically determined the block size using the following guidelines, where $L_{b}$ denotes the block size and $L_{i}$ denotes the image size:
$L_{b}=\left\{\begin{array}[]{ll}L_{i}/256&\text{if }L_{i}\geq 1024\text{ MiB},\\ L_{i}/64&\text{if 256 MiB }\leq L_{i}<1024\text{ MiB},\\ L_{i}/16&\text{if 16 MiB}\leq L_{i}<256\text{ MiB},\\ L_{i}&\text{otherwise.}\\ \end{array}\right.$ (1)
For smaller images, which can be quickly downloaded from a single source, we consolidate the image into a single block to minimize communication overhead while still applying peer scoring algorithms.
Peer Selection Mechanism. Upon determining the optimal block size, PeerSync proceeds to the critical phase of peer selection. This phase significantly impacts system efficiency, influencing both the download time and the load on network links, particularly in congested conditions. PeerSync’s selection algorithm evaluates peers based on three criteria: network position, network speed, and content popularity. Each peer is assigned a score within the range of [0,100][0,100][ 0 , 100 ], reflecting its suitability for selection, with higher scores correlating to a greater probability of being chosen for content retrieval. In the following, we refer to the device initiating the content request as the client.
Network Position and Speed. The first criterion in PeerSync’s peer selection is network position—whether a peer resides in the same LAN as the client. Local peers are prioritized, receiving the maximum score of 100. For peers outside the LAN, PeerSync employs a speed-based scoring algorithm, where higher network speeds result in higher scores. This strategy favors local peers when available, and otherwise selects the fastest peers. Let $s_{p}^{t}$ denote the speed from peer $p$ to the client at time step $t$, and $\bar{s}^{t}$ the average speed from all peers to the client at time step $t$. Additionally, PeerSync maintains a sliding window $\mathcal{W}_{p}$, of size $L$, for recording historical transmission speeds from peer $p$ to the client. Similarly, $\mathcal{W}$ records the global average speeds. The speed $s_{p}^{t}$ is updated as:
@@
and the global average speed $\bar{s}^{t}$ is calculated as:
@@
The network score $\text{net}_{p}^{t}$ of peer $p$ is then obtained as:
$\text{net}_{p}^{t}=s_{p}^{t}-\bar{s}^{t}.$ (4)
After re-scaling, $\text{net}_{p}^{t}$ is mapped to a value within [0,100][0,100][ 0 , 100 ].
Content Popularity. The popularity of an image layer affects its retrieval efficiency. Clients should avoid selecting peers storing rare images to optimize bandwidth usage. Let $\mathcal{P}^{t}$ denote the set of peers known to the client at time step $t$, and $\mathcal{I}_{p}^{t}$ the set of images possessed by peer $p$. We compute the popularity $\rho_{l}$ of layer $l$ as:
$\rho_{l}=\frac{\sum_{p\in\mathcal{P}^{t}}\sum_{i\in\mathcal{I}_{p}^{t}}\epsilon_{l}^{i}}{\sum_{p\in\mathcal{P}^{t}}\sum_{i\in\mathcal{I}_{p}^{t}}1},$ (5)
where $\epsilon_{l}^{i}=0$ if layer $l$ exists in image $i$, and 1 otherwise. The popularity score of peer $p$ is then given by:
$\text{pop}_{p}^{t}=100\times\left(1-\frac{\sum_{i\in\mathcal{I}_{p}^{t}}\sum_{l\in\mathcal{L}_{i}}e^{-\lambda\cdot\rho_{l}}}{\sum_{i\in\mathcal{I}_{p}^{t}}\sum_{l\in\mathcal{L}_{i}}1}\right).$ (6)
Custom Scoring and Softmax Distribution. Recognizing different user needs, PeerSync allows users to define custom scoring functions. The custom score of peer $p$ is denoted by $\text{cst}_{p}^{t}$. The final score (we also name it utility) of peer $p$, denoted by $U(p;t)$, is a weighted sum of network speed, content popularity, and the custom score:
@@
where $\alpha$, $\beta$, and $\gamma$ are weights. PeerSync uses a softmax function over the scores to compute the probability distribution for peer selection:
$\Pr\{p;t\}=\frac{\exp(U(p;t))}{\sum_{p^{\prime}\in\mathcal{P}}\exp(U(p^{\prime};t))}.$ (8)
This ensures peers with higher scores are more likely to be selected.
Theoretical Analysis. Let $p_{t}$ be the peer selected by PeerSync and $p_{t}^{*}$ the peer that maximizes utility for a given client at time $t$. The instantaneous regret for selecting peer $p_{t}$ is defined as:
$R_{t}=U(p_{t}^{*};t)-U(p_{t};t).$ (9)
Theorem 1.
PeerSync’s peer selection mechanism achieves sublinear regret in terms of the cumulative difference between the optimal utility and the utility obtained by PeerSync. Specifically, over $T$ rounds of peer selection for different clients and content requests, the cumulative regret is bounded by $O(\sqrt{T})$, i.e.,
$R(T)=\sum_{t=1}^{T}(U(p_{t}^{*};t)-U(p_{t};t))\leq O(\sqrt{T}).$ (10)
The proof is omitted.
Proof.
The regret bound follows from the softmax-based selection strategy employed by PeerSync, where the probability of selecting a peer is a function of its utility score. The softmax function, with a decaying temperature parameter $\tau_{t}$, ensures exploration early on and focuses on exploitation as $\tau_{t}\to 0$, converging to the optimal peer.
The expected instantaneous regret $R_{t}$ is bounded by:
$\mathbb{E}[R_{t}]\leq\frac{C}{\tau_{t}},$ (11)
where $C$ is a constant representing the maximum possible utility difference between peers. Using $\tau_{t}=\frac{\tau_{0}}{\sqrt{t}}$, the cumulative regret over $T$ rounds is bounded by
$R(T)\leq\sum_{t=1}^{T}\frac{C}{\sqrt{t}}=O(\sqrt{T}),$ (12)
which means that PeerSync’s peer selection mechanism has a sublinear regret. ∎
III-D Embedded Autonomous Tracker
Trackers serve as centralized services crucial for facilitating peer discovery within P2P networks [21]. Without a functional tracker, peer discovery times can increase dramatically due to the multi-hop iteration of DHT, leading to inefficiencies. However, the centralized architecture of trackers poses potential points of failure, such as hardware malfunctions, software crashes, network outages, or inadequate maintenance. Additionally, managing a separate tracker infrastructure introduces administrative overhead.
To address these challenges, PeerSync incorporates an embedded autonomous tracker system that dynamically initiates a leader election process when no active trackers are available within the swarm. This system is triggered when a node detects the absence of live trackers or fails to connect to an existing tracker. The autonomous tracker supports the coexistence of multiple trackers but activates its election mechanism when all trackers are unavailable, particularly in local swarm regions. The node that identifies this situation initiates the election process using the FloodMax algorithm. [^3] FloodMax is a well-known leader election protocol in distributed systems. See [28] for more details. This algorithm enables each node to broadcast its stability metrics, such as uptime, to neighboring nodes. To mitigate the high communication overhead inherent in the FloodMax algorithm for large-scale networks, PeerSync implements path-pruning techniques [33]. These techniques reduce unnecessary message traffic, especially across network boundaries, achieving near-linear message complexity. This optimization minimizes resource consumption while ensuring the timely election of a new tracker.
Integrated directly into PeerSync’s codebase, Embedded Autonomous Tracker eliminates the need for manual configuration, significantly reducing the risk of single points of failure. By enhancing resilience and adaptability, this feature ensures uninterrupted operations in dynamic and distributed environments.
III-E Cache Manager
Efficient storage management is essential for maintaining the stability of low-end edge devices, which often operate with constrained resources [29]. Over time, PeerSync accumulates residual images in its cache, consuming valuable disk space, even when the associated applications are no longer active. Furthermore, upper-level applications require spare storage for data-heavy tasks, such as storing raw sensor metrics and datasets. Edge devices, typically reliant on flash-based storage without advanced virtualization technologies like Ceph [45], face additional limitations. Consequently, PeerSync must maintain its storage footprint as small as possible while preserving performance.
PeerSync addresses these challenges with a dynamic cache cleaner that leverages usage history to manage storage. This component monitors image usage through the Docker Engine, gathering statistics to determine which images to remove when storage reaches critical levels (e.g., less than 10% free space or a user-defined threshold). Image removal is prioritized based on three factors: last access time, image size, and content popularity (we consider both local and global popularity).
PeerSync’s storage management algorithm extends the classic Least Recently Used (LRU) cache replacement strategy by incorporating an additional metric: cache miss cost. Unlike traditional LRU, which assumes uniform memory access cost across an address space [39, 40], PeerSync accounts for the non-trivial time differences in retrieving an image:
  • Images present within the same LAN incur minimal retrieval cost, making them candidates for removal if redundancy exists.
  • Images that are the only copy within a LAN but have multiple copies elsewhere are prioritized for removal based on the number of external replicas.
  • Sole known copies of images are given the lowest removal priority, ensuring critical data remains accessible unless absolutely necessary.
@@ III-E Cache Manager
This popularity-aware algorithm enables PeerSync to balance storage efficiency and performance, optimizing both retrieval times and disk utilization.
III-F Implementation Structure
PeerSync is designed with a message-passing architecture. Each component operates as an asynchronous task within the host system, with well-defined inter-component interfaces enabling communication and interaction. This approach provides both reliability and extensibility. In terms of reliability, isolated components ensure that failures do not affect the entire system. Any failed component can be quickly restarted to maintain continuous service. Extensibility is achieved through modular design, enabling easy updates to incorporate new features, adhere to evolving specifications, and support additional image formats.
PeerSync supports multiple peer discovery methods that function independently. This ensures Downloading Engine can immediately initiate fetching when an image is available from the P2P swarm, similar to BitTorrent. If P2P fetching is not feasible, PeerSync seamlessly falls back to direct HTTP downloading, ensuring uninterrupted operation.
@@ III-F Implementation Structure
langchain/langchain latest NLP 437.57 MB LLM application adaption framework[5]
@@
IV Evaluation

### v2

@@ III-C2 Downloading Engine
Table II: Layer size distribution of Docker Hub Top 100 images. Each percentage represents the proportion of layers that are smaller than the specified threshold.
Thres. (B) % << Thres. Thres. (B) % << Thres.
@@
256 Mi 97.19 605.73 Mi 100.00
The Downloading Engine is structured around a Request Dispatcher and a P2P Downloader.
Request Dispatcher
Upon receiving a layer digest from the API interface, the dispatcher first checks with the local Cache Manager (Sec. III-C4) and serves the content if locally present. Otherwise, a three-stage decision process is executed to retrieve the content from external sources. (i) It first queries the other nodes in the same LAN and fetches from the neighbor nodes. (ii) If not local, the dispatcher consults the manifest cache to obtain the layer’s size $L_{i}$ to determine if a P2P process shall be started. If $L_{i}$ is below a configurable threshold $\theta$ (e.g., 1 MiB), the request is routed directly to the upstream registry. This is justified by the data in Table II, which shows that a policy of direct fetch for small layers can handle a large fraction of requests with minimal latency. (iii) For layers larger than $\theta$, the dispatcher initiates a P2P discovery process. However, this process is bounded by a dynamic timeout $\tau$, defined as the 95th percentile of recent round-trip times (RTTs) over a 10-second sliding window. If no suitable peer is found within $\tau$, the system falls back to the upstream registry to avoid indefinite user wait times. This logic can be viewed as a state machine (Fig. 4), but is now implemented as a streamlined, asynchronous workflow to reduce overhead.
Popularity- and Network-Aware P2P Downloader
PeerSync segments each image layer into fixed-size blocks to enable concurrent downloads from multiple peers. Upon receiving a download request from the Request Dispatcher, the P2P Downloader initiates a five-stage workflow (Fig. 5): (i) select a batch of blocks for the current cycle; (ii) sample candidate peers based on dynamically updated scores maintained by the Peer Aggregator; (iii) assign each block to the highest-scoring peer; (iv) issue download requests; and (v) validate received blocks via cryptographic hash. During this phase, PeerSync re-queues failed blocks or caching and delivers valid ones.
@@
Determining the block size. Block size critically balances parallelism and overhead. Excessively large blocks limit concurrency; overly small ones inflate Merkle tree depth and hash computation costs. Empirical evaluation using an 8194.5 MiB image in a 10 Gbps LAN (Table III) shows optimal performance at 16-128 MiB. Accordingly, PeerSync sets block size $L_{b}$ adaptively based on image size $L_{i}$:
[…]
For tiny images whose size $L_{i}<16$ MiB, a single block is used to minimize coordination overhead.
Peer selection rules. Peer selection combines three criteria into a unified utility score $U(p;t)\in[0,100]$: network position content popularity, and customized logic.
  1. 1.
    Network position. Peer selection uses a network-aware score that favors peers with higher effective throughput. LAN-local peers receive the maximum score of 100. For remote peers, the score is based on observed download speed, not low-level metrics like packet loss, as congestion control is handled by the OS kernel [34, 35, 36, 37]; using throughput avoids interfering with kernel mechanisms and better reflects end-to-end performance. Each peer $p$ maintains a sliding window $\mathcal{W}_{p}$ of past speeds $\{s_{p}^{t^{\prime}}\}$. The current speed estimate $s_{p}^{t}$ is an exponentially weighted average that prioritizes recent samples:
@@
    where $L$ is the current logical time. A global baseline $\bar{s}^{t}$ is computed similarly over all recent transfers:
@@
    The raw network advantage is $\text{net}_{p}^{t}=s_{p}^{t}-\bar{s}^{t}$, which is then linearly rescaled to $[0,100]$ (with negatives clamped to the minimum known score) to yield the final network score. This relative scoring adapts to heterogeneous and dynamic network conditions.
  2. 2.
    Content popularity. Content popularity discourages reliance on peers holding rare layers. Let $\rho_{l}\in[0,1]$ denote the fraction of known image instances containing layer $l$:
    $\rho_{l}=\frac{\sum_{p\in\mathcal{P}^{t}}\sum_{i\in\mathcal{I}_{p}^{t}}\epsilon_{l}^{i}}{\sum_{p\in\mathcal{P}^{t}}\sum_{i\in\mathcal{I}_{p}^{t}}1},\quad\epsilon_{l}^{i}=\begin{cases}1&\text{if }l\in i,\\ 0&\text{otherwise}.\end{cases}$ (4)
    Peers periodically exchange layer sets via differential updates; unchanged states are acknowledged with sequence numbers to minimize overhead. The popularity score for peer $p$ is then:
    $\text{pop}_{p}^{t}=100\times\left(1-\frac{\sum_{i\in\mathcal{I}_{p}^{t}}\sum_{l\in\mathcal{L}_{i}}e^{-\lambda\cdot\rho_{l}}}{\sum_{i\in\mathcal{I}_{p}^{t}}\sum_{l\in\mathcal{L}_{i}}1}\right),$ (5)
    which down-weights peers storing rarer content. The rationale is to preserve the bandwidth of these critical peers, ensuring they remain available to serve the rare content that only they possess, thus enhancing the overall availability of all layers in the swarm.
  3. 3.
    Extensibility via custom scoring. To accommodate the diverse and heterogeneous nature of edge deployments, PeerSync’s scoring function is designed with an extensibility hook for domain-specific policies, represented by a custom score $\text{cst}_{p}^{t}$. This architectural choice allows administrators to inject logic tailored to their unique operational constraints without modifying the core system. For instance, in a wireless edge environment, an administrator may wish to deprioritize peers with weak signal strength to conserve their battery and bandwidth. This can be achieved by defining the custom score $\text{cst}_{p}^{t}$ as a function of the peer’s real-time signal strength, thereby penalizing nodes with poor connectivity in the selection process.
The total utility is a weighted sum:
@@
with $\alpha+\beta+\gamma=1$. In the default configuration, $\alpha=\beta=0.5$, and custom scoring is disabled. The three weights may be adjusted accordingly to achieve a balance between network optimization and resilience. A high $\alpha$ value indicates a strong preference for nearby peers, while $\beta$ strengthens availability. $\gamma$ is configured along with the custom score for more flexibility. Eventually, peer selection uses a softmax sampling over the utility scores $U(p;t)$, favoring high-utility peers while preserving space for exploration.
III-C3 Embedded Autonomous Tracker
In P2P networks, trackers play a central role in accelerating peer discovery [17]. Without an active tracker, nodes must fall back to multi-hop DHT lookups, which significantly increase discovery latency and degrade download performance. However, traditional trackers introduce a single point of failure, vulnerable to crashes, network partitions, or misconfiguration, and require separate deployment and maintenance.
To eliminate this dependency, PeerSync embeds an autonomous tracker module that self-activates when no live trackers are reachable. Specifically, if a node fails to contact any known tracker during bootstrap or periodic health checks, it initiates a leader election within its local network partition using the FloodMax algorithm [38]. In this protocol, each candidate broadcasts a stability score (based on metrics including uptime and known neighbors) and propagates the highest score observed so far. The node with the globally maximal score becomes the new tracker. While FloodMax ensures correctness, its naive implementation incurs $O(n^{2})$ message complexity in dense networks. To address this, PeerSync applies path-pruning techniques [39] that suppress redundant message forwarding across already-explored paths and avoid cross-subnet flooding unless necessary. This reduces communication overhead to near-linear complexity while preserving convergence guarantees.
Since the tracker logic is fully integrated into PeerSync’s runtime, no external configuration or manual intervention is required. This design not only removes operational overhead but also enhances system resilience, which can ensure continuous peer discovery even in transiently partitioned or infrastructure-free edge environments.
III-C4 Cache Manager
Efficient storage management is critical for low-end edge devices, which typically operate under severe resource constraints and rely on flash-based storage without advanced virtualization layers like Ceph [40, 41]. Over time, PeerSync accumulates cached images, even after their associated applications have terminated, consuming disk space needed for data-intensive tasks such as storing sensor logs or datasets. To preserve system stability while maintaining performance, PeerSync must minimize its storage footprint without compromising image availability. To this end, PeerSync implements a dynamic cache cleaner that monitors image usage via the container runtime and triggers eviction when free space falls below a threshold (e.g., 10% or a user-defined limit). Eviction decisions are guided by three factors: (i) last access time, (ii) image size, and (iii) content popularity, considering both local presence and global replication.
The underlying algorithm extends the classic Least Recently Used (LRU) policy by incorporating cache miss cost, which reflects the non-uniform cost of retrieving an image after eviction [42, 43]. Unlike RAM-based caching, in which access latency is uniform, image retrieval in PeerSync varies significantly by network context:
  • Images with multiple local replicas (within the same LAN) incur near-zero miss cost and are safe to evict.
  • Images that are the sole local copy are assigned a miss cost inversely proportional to the number of remote replicas.
  • Sole known copies are assigned the highest retention priority and evicted only under extreme storage pressure.
@@ III-C4 Cache Manager
By jointly optimizing recency, space, and retrieval cost, this popularity-aware strategy enables PeerSync to balance disk utilization and performance on resource-constrained edge nodes.
@@
langchain/langchain latest NLP 437.57 MB LLM application adaptation framework[46]
@@
III-D Implementation Concerns
PeerSync adopts a message-passing architecture in which each functional component runs as an independent asynchronous task. Components communicate exclusively through well-defined, typed message interfaces, decoupling internal logic from inter-component coordination. A crash or stall in one module (e.g., due to a malformed image manifest or network timeout) does not propagate to others. The supervisor process can restart the failed component within milliseconds, ensuring continuous operation without full-system disruption. In addition, new peer discovery protocols, storage backends, or scoring heuristics can be integrated by implementing standardized interfaces without modifying core logics. This facilitates rapid adaptation to upcoming container formats or edge-specific constraints.
Peer discovery leverages multiple orthogonal mechanisms, centralized trackers, Kademlia-based DHT, and LAN-local multicast, all operating concurrently. The Downloading Engine monitors availability signals from these sources and initiates block fetching as soon as any peer in the P2P swarm reports possession of the requested layer, mimicking BitTorrent’s opportunistic pull model. If no P2P sources are reachable within a configurable timeout (or if the content is globally unique), PeerSync transparently falls back to direct HTTP(S) retrieval from the upstream registry. This hybrid strategy guarantees liveness while maximizing bandwidth utilization whenever collaborative sources exist.
IV Adaptive Peer Selection and System-Level Guarantees
Peer selection in PeerSync operates in a non-stationary, partially observable environment where performance and content distribution vary constantly. Rather than pursuing unattainable optimality guarantees against an ill-defined oracle, we analyze the mechanism through three analytically tractable lenses: (i) convergence of utility estimation under exponential smoothing, (ii) probabilistic protection of rare-content peers, and (iii) approximation quality of aggregate throughput maximization. Together, these properties explain why the heuristic consistently achieves high speedup while preserving swarm health.
IV-A Convergence of Network Utility Estimation
The network score $\text{net}_{p}^{t}$ is derived from an exponentially weighted moving average (EWMA) of observed download speeds (see (2)), where $\tau_{s}>0$ is the smoothing time constant (implicit in your original notation via $L-t^{\prime}$). Let $B_{p}(t)$ denote the true instantaneous throughput achievable from peer $p$ at time $t$, and assume $|s_{p}^{t^{\prime}}-B_{p}(t^{\prime})|\leq\sigma$ for all $t^{\prime}$ (bounded measurement noise). Then, under mild Lipschitz continuity of $B_{p}(\cdot)$ (i.e., $|B_{p}(t)-B_{p}(t-1)|\leq\delta$), the estimation error satisfies
$|s_{p}^{t}-B_{p}(t)|\leq\sigma+\delta\cdot\tau_{s}.$ (7)
Thus, by choosing $\tau_{s}$ appropriately (e.g., $\tau_{s}=2$ rounds in our implementation), the EWMA estimator tracks the true throughput within a bounded error envelope. Since $\text{net}_{p}^{t}$ is a monotonic rescaling of $s_{p}^{t}-\bar{s}^{t}$, it inherits this stability, ensuring that transient anomalies do not dominate peer ranking.
IV-B Probabilistic Protection of Rare-Content Peers
Let $\mathcal{R}_{t}=\{l:\rho_{l}<\rho_{\min}\}$ denote the set of rare layers at time $t$, and define the rarity exposure of peer $p$ as
$\eta_{p}^{t}=\frac{1}{|\mathcal{L}_{p}^{t}|}\sum_{l\in\mathcal{L}_{p}^{t}}\mathbb{I}[l\in\mathcal{R}_{t}],$ (8)
where $\mathcal{L}_{p}^{t}$ is the set of layers held by $p$ and $\mathbb{I}[\cdot]$ is the indicator function. The popularity score can be upper-bounded as
$\text{pop}_{p}^{t}\leq 100\cdot(1-e^{-\lambda\rho_{\min}})\cdot(1-\eta_{p}^{t})+100\cdot(1-e^{-\lambda})\cdot\eta_{p}^{t}.$ (9)
Consequently, if $\eta_{p}^{t}\geq\eta_{0}$ (i.e., peer $p$ holds many rare layers), then $\text{pop}_{p}^{t}\leq\bar{u}_{\text{rare}}<100$. Under softmax selection with temperature $\tau$, the probability of selecting such a peer is upper-bounded by
$\Pr\{p_{t}=p\}$ $\leq\frac{\exp\big((\alpha\cdot 100+\beta\cdot\bar{u}_{\text{rare}})/\tau\big)}{\exp\big((\alpha+\beta)\cdot 100/\tau\big)}$
$=\exp\left(-\frac{\beta(100-\bar{u}_{\text{rare}})}{\tau}\right).$ (10)
This exponential suppression ensures that peers critical for rare content are selected infrequently, thereby conserving their bandwidth. In steady state, this mechanism bounds the expected depletion rate of rare layers, enhancing long-term availability.
IV-C Throughput Approximation via Utility Maximization
Let $\mathcal{S}_{t}\subseteq\mathcal{P}$ be the set of $k$ peers selected for concurrent download at time $t$. Assume block-level parallelism and negligible coordination overhead, so the total bandwidth is
$B_{\text{total}}(t)=B_{0}+\sum_{p\in\mathcal{S}_{t}}B_{p}(t).$ (11)
Since $\text{net}_{p}^{t}$ is a linear transformation of $s_{p}^{t}$, and $s_{p}^{t}$ approximates $B_{p}(t)$ within error $\epsilon=\sigma+\delta\tau_{s}$, we have
$\left|\sum_{p\in\mathcal{S}_{t}}\text{net}_{p}^{t}-\sum_{p\in\mathcal{S}_{t}}B_{p}(t)\right|\leq k\cdot C\cdot\epsilon,$ (12)
for some constant $C$ from the rescaling. Because $\text{pop}_{p}^{t}\in[0,100]$ is independent of instantaneous bandwidth, the total utility sum satisfies
$\sum_{p\in\mathcal{S}_{t}}U(p;t)=\alpha\sum_{p\in\mathcal{S}_{t}}\text{net}_{p}^{t}+\beta\sum_{p\in\mathcal{S}_{t}}\text{pop}_{p}^{t}+\gamma\sum_{p\in\mathcal{S}_{t}}\text{cst}_{p}^{t}.$ (13)
Thus, maximizing $\sum_{p\in\mathcal{S}_{t}}U(p;t)$ approximately maximizes $\sum_{p\in\mathcal{S}_{t}}B_{p}(t)$ up to an additive error of $O(k\epsilon)$. In practice, since $\alpha=\beta=0.5$, the scheduler jointly optimizes for speed and resilience, achieving a Pareto-efficient trade-off.
V Evaluation